OpenStudy (anonymous):

A Sample of neon is collected at 2.7 atm and 12.0 C. It has a volume of 2.25 L. What would be the volume of this gas at STP. It says that i have to use the combined gas law,but i do not know if i should plugs these number in for P1V1/T1 or P2V2/T2 help thank you

OpenStudy (anonymous):

P1V1/T1 = P2V2/T2 252cm3

OpenStudy (cyber405):

This equation is for ideal gases...if we suppose neon as a ideal gas then we can use it.. P1V1/T1=P2V2/T2 P1=2.7 atm P2=1 atm V1=2.25 L V2=??? L T1=12+273=285 K T2=273

OpenStudy (cyber405):

According to new condition for neon (which is STP) we know that the volume of gas should increase... Is that helpful...???